A person plans to invest three times as much in an account at 7% interest as in an account at 9% annual interest. how much should the person invest in each account to earn a total of $300 in one year.
The person should invest $---- at 7% annual interest and $----- at 9% annual interest.

Answer by mananth(16946)

at 9%-------------------x
at 7%-------------------3x
Interest
9%x+7%3x=300
multiply by 100
9x+7x=30000
16x=30000
/16
x=1875
$1875 @9%
$5625@7%